Suppose Rosmary O’Leary is deciding how to dispose of some hazardous waste. She can contain it safely at a cost of $175 or bury it in the local landfill. If she chooses the second option, in ten years’ time the stuff will have seeped out enough to permanently ruin the vineyard of Emelle Alabama, who lives next-door to the landfill. The vineyard has a value to her in ten years of $450 (she’d sell it then for $450).

a. What is the present value of $450 in ten years at a discount rate of 10%?

b. Is it socially efficient to bury the waste in the landfill?
Answer #1

ANSWER:

A) Pv = 450(p/f,i,n) = 450(p/f,10%,10) = 450 * 0.3855 = 173.475

B) Since the present value of $450 is less then the cost of burying in the local landfill that is 173.475 < 175 , therefore it is not socially efficient to bury the waste in the local landfill.